an object of mass m is dropped from a height h.what is the kinetic energy when it touches the ground?"
1
Expert's answer
2013-01-10T09:37:06-0500
T+U=E=const - The law of conservation of energy
T - kinetic energy
U - potential energy
E - total energy
when the object is dropped:
E=0+mgh
m - mass oа the body
h - height
when it touches the ground
E=T+0
T+0=0+mgh
T=mgh